OpenStudy (anonymous):

int_{}^{?}e ^{4t}sin3t

OpenStudy (anonymous):

integrate \[\int\limits_{}^{?}e ^{4t}\sin3t\]

OpenStudy (anonymous):

\[\int\limits e ^{4t} \sin3t dt ?\]

OpenStudy (anonymous):

can be solved by integration by parts

OpenStudy (anonymous):

\[e ^{4t} -\cos3t/3 -\int\limits -\cos3t/3.e ^{4t} 4 dt\]

OpenStudy (anonymous):

\[\cos3t.e ^{4t} /3 +(4/3)\int\limits e ^{4t} \cos 3t d t\]

OpenStudy (anonymous):

oh i missed the negative sign -(cos3t.e4t)/3+(4/3)∫e4tcos3tdt

OpenStudy (anonymous):

yea its by using the formula

OpenStudy (anonymous):

Yeah, you will never finish integrating but you will get a multiple of the original integral eventually...so you move it to the other side of the equation and combine with the original integral and, shazam!, you are done.